OpenStudy (anonymous):

Write the polar equation r = cos(4*theta) in the parametric form x = x(t), y = y(t)

OpenStudy (anonymous):

multiply both sides by r r^2 = 4 cos (4theta) r^2 = 8cos^4 ( theta) - 8 cos^2(theta) + 1

OpenStudy (anonymous):

x^2+y^2 = 8 ( x/r)^4 - 8(x/r)^2 + 1